# Multi-Modal Attention Networks with Uncertainty Quantification for Automated Concrete Bridge Deck Delamination Detection

### v1

@@ Abstract
Deteriorating civil infrastructure requires automated inspection techniques that overcome limitations of traditional visual assessment. While Ground Penetrating Radar and Infrared Thermography enable subsurface defect detection, single-modal approaches face complementary constraints—radar struggles with moisture interference and shallow defects, while thermography exhibits weather dependency and limited penetration depth. This paper presents a multi-modal attention network that fuses radar temporal patterns with thermal spatial signatures for automated bridge deck delamination detection. Our architecture introduces temporal attention for radar signal processing, spatial attention for thermal feature extraction, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns invisible to individual sensors. We incorporate uncertainty quantification through Monte Carlo dropout and learned variance estimation, decomposing predictive uncertainty into epistemic and aleatoric components essential for safety-critical decisions. Experiments on five bridge datasets reveal that on balanced to moderately imbalanced data, our approach substantially outperforms baselines in accuracy and area under curve—representing meaningful improvements over single-modal approaches and concatenation-based fusion. Ablation studies demonstrate that cross-modal attention provides critical gains beyond within-modality attention alone, while multi-head mechanisms achieve improved calibration over single-head alternatives. Uncertainty quantification reduces calibration error, enabling selective prediction by rejecting uncertain cases for human review. However, under extreme class imbalance, attention mechanisms demonstrate vulnerability to majority class collapse. These findings provide actionable guidance: attention-based architecture performs well across typical inspection scenarios, while extreme imbalance requires specialized techniques. Our system maintains deployment efficiency with compact parameterization, enabling real-time bridge inspection with characterized capabilities and limitations.
Keywords: Bridge deck inspection; Multi-modal fusion; Attention mechanisms; Uncertainty quantification; Infrastructure monitoring
@@ 1  Introduction
Aging infrastructure presents a critical challenge globally, with the American Society of Civil Engineers estimating a \$2.6 trillion investment gap for U.S. infrastructure through 2029 [2]. Bridge deck deterioration alone affects over 231,000 structures nationwide, with delamination representing the primary failure mode in concrete bridge decks [7]. Traditional inspection methods fail to provide the comprehensive, objective assessment needed for effective maintenance planning and risk management.
Visual inspection remains the predominant assessment method for concrete infrastructure. However, surface-based evaluation cannot detect subsurface delaminations, chloride contamination, or early-stage deterioration that compromises structural integrity [1]. Manual inspection introduces significant subjectivity and variability between inspectors. Weather dependencies, accessibility constraints, and safety risks further limit inspection effectiveness. These limitations necessitate objective, comprehensive evaluation techniques capable of detecting both surface and subsurface defects with quantifiable confidence levels.
@@
The complementary nature of these sensing modalities—where GPR provides depth penetration that IRT lacks, while IRT captures shallow defects and surface detail that challenge GPR—motivates multi-modal fusion approaches. However, realizing this complementarity requires fusion mechanisms capable of learning complex inter-modal relationships rather than simply combining features. Conventional fusion approaches using concatenation or weighted averaging treat modalities as independent information sources, failing to capture the sophisticated correspondences between electromagnetic reflection patterns and thermal signatures that jointly indicate delamination characteristics. Recent research demonstrates that attention mechanisms fundamentally outperform traditional fusion strategies by 10–30% across infrastructure applications [16, 27]. Attention mechanisms provide selective feature weighting that dynamically assigns importance based on context, cross-modal interaction modeling that captures relationships between sensing modalities, and global context modeling through self-attention that enables long-range dependency capture across spatial and temporal dimensions [25, 23].
Vision Transformers have achieved strong performance in infrastructure defect detection, with recent studies reporting accuracy rates exceeding 90% while outperforming CNN-based approaches [14]. However, critical knowledge gaps remain regarding optimal fusion strategies for time-series sensor data, uncertainty quantification for safety-critical deployment, and crucially, understanding under what conditions attention-based architectures provide advantages over simpler alternatives [33, 22]. This final question proves particularly important for practical deployment: bridge inspection datasets exhibit diverse class distributions ranging from relatively balanced to severely imbalanced, and the relationship between dataset characteristics and optimal modeling approach remains poorly understood.
@@
This work addresses these gaps by introducing a multi-modal attention network that fuses GPR and IRT data for automated bridge deck delamination detection with integrated uncertainty quantification. Beyond demonstrating strong performance, we conduct systematic investigation into when and why attention-based fusion succeeds, providing actionable guidance for practitioners. The proposed architecture combines temporal attention for GPR signal processing, spatial attention for IRT thermal feature extraction, and multi-head cross-modal fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities. This architectural design integrates comprehensive uncertainty quantification that decomposes predictive uncertainty into epistemic and aleatoric components through Monte Carlo dropout and learned variance estimation, enabling selective prediction strategies that achieve 93% accuracy by rejecting uncertain cases for human review.
@@
We validate the proposed approach through systematic experimental investigation across five bridge datasets from the SDNET2021 benchmark [12], revealing that attention-based fusion substantially outperforms baselines on balanced to moderately imbalanced datasets while exhibiting specific vulnerability under extreme class imbalance. These findings provide practical deployment guidance rather than claims of universal superiority. Comprehensive ablation studies further demonstrate that cross-modal attention provides critical performance gains beyond within-modality attention alone, with multi-head mechanisms achieving substantially improved calibration over single-head alternatives.
@@
The cross-modal attention modules learn optimal feature interaction patterns between electromagnetic and thermal signatures, while Bayesian uncertainty quantification provides confidence intervals for defect predictions enabling risk-informed maintenance decisions [8]. Experimental validation demonstrates that on four of five bridge datasets representing typical inspection scenarios, the architecture achieves 10–25 percentage point improvements in accuracy and AUC compared to single-modal approaches and simple fusion baselines, with particularly strong results on the Park River Median dataset (96.6% accuracy, 99.8% AUC). The investigation also characterizes boundary conditions where architectural simplification or specialized imbalance handling may be preferable, providing the nuanced understanding essential for reliable deployment in safety-critical infrastructure applications.
@@ 2  Literature Review
The deterioration of civil infrastructure presents a critical challenge requiring accurate detection of subsurface defects in concrete structures. Traditional single-modal non-destructive evaluation (NDE) approaches have demonstrated fundamental limitations in comprehensive defect characterization. Ground penetrating radar (GPR) alone, while effective for detecting subsurface anomalies, struggles with moisture-related artifacts and requires expert interpretation [28, 6]. Infrared thermography (IRT) provides excellent surface temperature mapping but lacks depth penetration capabilities, limiting its effectiveness for deep delamination detection [10]. Impact echo testing offers precise depth information but requires dense sampling for comprehensive coverage, making it time-intensive for large-scale assessments [32]. Research has consistently shown that individual NDE methods achieve accuracy rates between 75–85% for delamination detection, with significant variations based on environmental conditions and defect characteristics [9]. These limitations have driven the development of multi-modal approaches that leverage complementary information from multiple sensors to overcome individual modality constraints.
Initial attempts at multi-modal fusion employed simple concatenation strategies, where features from different NDE modalities were combined at either early or late fusion stages. Omar and Nehdi [20] demonstrated that combining GPR and IRT data through weighted averaging achieved 5–8% accuracy improvements over single-modal approaches. However, these early methods suffered from several critical limitations: loss of modality-specific features during concatenation, inability to handle varying data dimensions effectively, and lack of adaptive weighting mechanisms for different reliability conditions. The introduction of the SDNET2021 dataset marked a significant milestone, providing 663,102 annotated GPR signals, 1,936 Impact Echo signals, and 4,580,680 IRT pixels from five real in-service bridge decks [12]. The dataset’s three-class annotation scheme—intact, shallow delamination above reinforcement, and deep delamination below reinforcement—aligns with standard bridge deck repair protocols. Critically for systematic investigation, the five bridge datasets within SDNET2021 exhibit substantially different class distributions, ranging from relatively balanced to severely imbalanced with intact regions exceeding 90% of samples. This natural variation provides an opportunity to investigate how model performance varies with dataset characteristics—an analysis not conducted in prior work using this benchmark.
The first generation of deep learning-based fusion architectures addressed the concatenation problem through hierarchical feature extraction. Pozzer et al. [24] proposed a dual-stream CNN architecture that processed GPR and IRT data through separate encoders before fusion, achieving 87.3% classification accuracy on concrete delamination tasks. This approach preserved modality-specific features through dedicated processing pathways while enabling learned fusion at multiple network depths. Multi-level feature fusion strategies emerged as a solution to capture both low-level details and high-level semantic information. Liu et al. [18] demonstrated that fusing features at early, intermediate, and late stages improved F1-scores by 12–15% compared to single-stage fusion. Their architecture employed skip connections to preserve fine-grained features while allowing the network to learn optimal combination strategies at each level. However, these approaches still relied on fixed fusion weights, revealing the need for mechanisms that could dynamically adjust fusion strategies based on modality reliability and data quality.
The introduction of attention mechanisms enabled selective feature focusing and dynamic weight adjustment, representing a significant advance over fixed-weight fusion. Spatial attention mechanisms proved particularly effective for IRT data, with channel and spatial attention combinations achieving 10–12% AUC improvements over baseline CNN approaches [29]. The spatial attention formulation enabled networks to focus on thermally anomalous regions indicative of subsurface defects while suppressing background noise through a combination of average and max pooling operations followed by convolutional filtering. Temporal attention mechanisms similarly transformed GPR signal processing by identifying critical reflection patterns in time-series data. The Temporal-Spatial Cross-Attention Networks (TSCA-Net) architecture [4] combined LSTM with Transformers to capture temporal dependencies, achieving 92.6% accuracy using only 5% of available training labels, while reducing computational requirements by 40%. The Multi-modal Fusion Network (M²FNet) combined CNN-based local feature extraction with Transformer global characterization [17], employing cross-attention mechanisms to fuse raw signal data with B-scan and top-scan GPR images and achieving 95% AUC scores. Vision Transformers adapted for defect detection further advanced the field, with the Multi-feature Vision Transformer (M-VIT-DDQ) achieving 98.31% defect recognition accuracy [33], while the Dynamic Sparse Attention Transformer (DSAT) architecture [22] introduced hierarchical feature fusion with bidirectional dense connections, reducing computational complexity by 60% while maintaining accuracy. These results demonstrate strong average performance for attention-based architectures; however, existing studies predominantly evaluate methods on datasets with relatively balanced class distributions or apply class balancing techniques prior to evaluation. Whether the multi-head attention layers that prove effective on balanced data maintain their advantages when minority class samples provide limited gradient signal remains an open question with significant practical implications.
Infrastructure assessment demands not only accurate predictions but also calibrated confidence estimates for safety-critical decision-making. Monte Carlo Dropout emerged as the predominant uncertainty quantification method, with research demonstrating that 15–25 forward passes provide reliable epistemic uncertainty estimates [8]. Yang et al. [31] applied this framework to infrastructure monitoring, showing that total predictive uncertainty could be effectively decomposed into aleatoric uncertainty arising from data noise and epistemic uncertainty from model limitations. The Bayesian Boundary-Aware Convolutional Network (B-BACN) advanced uncertainty-aware crack detection, achieving 84.7% F1-score with 93.0% accuracy while providing decomposed uncertainty estimates [21]. The approach revealed that mean class softmax variance and entropy strongly correlate with misclassifications, enabling uncertainty-triggered human intervention. Copula-based uncertainty quantification addressed multi-sensor fusion scenarios, decoupling univariate marginal probability density functions from dependence structure [15]. This statistical approach proved robust to sensor failures, maintaining 85% prediction accuracy even with 40% sensor dropout, while temperature scaling has emerged as the standard post-hoc calibration technique achieving 15–30% reduction in Expected Calibration Error.
The severe class imbalance inherent in infrastructure data—typically presenting as 77% intact regions, 17% shallow delamination, and 6% deep delamination—has motivated extensive research into specialized training strategies. Focal Tversky Loss emerged as a robust approach, outperforming 11 alternative loss functions across multiple benchmarks [13]. The focal component controls the focus on hard examples, achieving 23% improvement in minority class detection compared to standard cross-entropy loss. Weighted Binary Cross-Entropy with inverse frequency weighting provided a simpler alternative with comparable performance [26]. The Multi-Modal Industrial Surface Defect Detection (MISDD) framework introduced cross-modal prompt learning to handle missing sensor scenarios [30], maintaining 90% of full-modal performance even with 50% modality dropout. However, existing class imbalance research focuses predominantly on loss function modifications and sampling strategies, implicitly assuming that the underlying architecture remains equally effective regardless of class distribution. Whether attention mechanisms—with their increased parameter capacity and learned weighting schemes—exhibit different sensitivity to class imbalance compared to simpler CNN architectures has not been systematically investigated. This gap is particularly relevant given that attention layers learn to weight features based on training data statistics; when minority classes contribute fewer than 2% of training samples, the attention heads may receive insufficient gradient signal to develop specialized minority class detectors, potentially causing representational collapse toward majority class patterns.
Edge computing optimization has become critical for practical deployment, with model quantization techniques achieving real-time processing on resource-constrained devices [19]. TensorRT Float16 quantization reduced model size by 50% while maintaining 97% of original accuracy, and integer quantization achieved fourfold speedup with acceptable 3% accuracy degradation, enabling deployment on edge devices with 26–28 fps processing rates. Knowledge distillation from large transformer models to efficient CNN architectures provided another optimization path, with student models achieving 95% of teacher performance at 10% of computational cost [3].
The evolution from simple concatenation to attention-based architectures with uncertainty quantification represents substantial progress in automated infrastructure assessment, with studies consistently reporting accuracy improvements of 10–30% over traditional fusion approaches. However, this body of work leaves critical questions unanswered. Existing evaluations predominantly report average performance metrics without systematically investigating how results vary across datasets with different characteristics, leaving practitioners without guidance for selecting appropriate methods for their specific inspection scenarios. While attention mechanisms demonstrate clear advantages on benchmark evaluations, studies have not examined whether the architectural properties that enable these advantages—learned feature weighting, multi-head specialization, cross-modal interaction modeling—might become liabilities under certain data conditions. The increased model capacity that benefits learning on balanced data could potentially enable overfitting to majority class patterns when minority classes are severely underrepresented. Furthermore, uncertainty quantification research has developed largely independently from attention mechanism research, with limited investigation of how these capabilities interact and whether attention-based architectures produce better-calibrated uncertainty estimates than simpler alternatives. These gaps motivate our systematic investigation combining cross-modal attention fusion with comprehensive uncertainty quantification, evaluated across five bridge datasets exhibiting diverse class distributions. Rather than claiming universal superiority, we aim to characterize when attention-based fusion provides clear advantages, identify conditions under which simpler approaches may be preferable, and provide actionable guidance for practitioners deploying automated inspection systems across real-world infrastructure networks.

### v3

# Multi-Sensor Attention Networks for Automated Subsurface Delamination Detection in Concrete Bridge Decks

@@ Abstract
Subsurface delaminations in concrete bridge decks remain undetectable through conventional visual inspection, necessitating automated non-destructive evaluation methods. This work introduces a deep learning framework that integrates Ground Penetrating Radar (GPR) and Infrared Thermography (IRT) through hierarchical attention mechanisms. Our architecture employs temporal self-attention to process GPR electromagnetic signals, spatial attention to analyze thermal imagery, and cross-modal attention with learnable embeddings to model inter-sensor correspondences. We integrate Monte Carlo dropout-based uncertainty quantification, decomposing prediction confidence into model uncertainty and data-driven uncertainty components. Testing across five real-world bridge datasets from the SDNET2021 benchmark reveals that our approach delivers substantial performance gains over single-sensor and concatenation-based baselines when applied to balanced or moderately imbalanced data distributions. Comprehensive ablation analysis confirms that cross-modal attention mechanisms contribute meaningful improvements beyond unimodal attention alone. Critically, we identify and characterize specific failure modes: under extreme class imbalance, attention-based architectures demonstrate susceptibility to majority class bias, indicating scenarios where simpler architectural choices may prove more robust. Our findings equip practitioners with empirically-grounded criteria for selecting appropriate fusion strategies based on dataset characteristics, rather than promoting universal architectural superiority.
Keywords: Non-destructive evaluation; Multi-modal deep learning; Infrastructure monitoring; Sensor fusion; Concrete delamination
@@ 1  Introduction
The structural integrity of civil infrastructure presents mounting challenges as existing systems age beyond their designed service life. Current assessments indicate a \$2.6 trillion infrastructure investment deficit through 2029 in the United States alone [2], with bridge deck deterioration representing a particularly acute concern affecting over 231,000 structures nationwide [5]. Delamination—the separation of concrete layers from embedded reinforcement—constitutes the predominant degradation mechanism in reinforced concrete bridge decks, yet remains invisible to surface-based evaluation methods.
Contemporary bridge inspection practice relies heavily on visual assessment techniques that fundamentally cannot detect subsurface anomalies. This limitation extends beyond simple detection failure: visual methods introduce substantial inter-inspector variability, depend critically on weather conditions and site accessibility, and present significant safety hazards [1]. The resulting inspection paradigm fails to provide the objective, comprehensive data required for evidence-based maintenance planning and infrastructure risk management. These constraints have driven sustained research into non-destructive evaluation (NDE) technologies capable of characterizing both surface and subsurface deterioration with quantifiable reliability.
@@
This complementary coverage—GPR providing depth penetration where IRT struggles, IRT capturing surface detail where GPR fails—creates strong theoretical motivation for sensor fusion. However, effectively exploiting this complementarity requires fusion mechanisms capable of learning complex inter-modal relationships rather than merely combining feature vectors. Traditional fusion approaches using concatenation or weighted summation treat sensing modalities as independent information sources, failing to model the sophisticated correspondences between electromagnetic reflection patterns and thermal anomaly signatures that jointly characterize delamination. Recent infrastructure assessment research demonstrates that attention mechanisms achieve 10–30% performance improvements over conventional fusion by enabling selective feature weighting, cross-modal interaction modeling, and global context integration through self-attention [10, 14]. Vision Transformer architectures have demonstrated strong infrastructure defect detection performance exceeding 90% accuracy [9], though critical gaps remain regarding optimal fusion strategies for time-series sensor data, uncertainty quantification requirements for safety-critical deployment, and—crucially—understanding the boundary conditions under which attention-based architectures provide advantages over simpler alternatives [16, 11].
This final question proves particularly important for practical deployment. Bridge inspection datasets exhibit substantial diversity in class distributions ranging from relatively balanced to severely skewed, yet the relationship between dataset characteristics and optimal architectural choice remains poorly characterized. Understanding not only when attention-based fusion excels but also when it may prove inferior to simpler approaches is essential for reliable deployment in safety-critical infrastructure applications.
@@
This work addresses these gaps through development and systematic evaluation of a multi-modal attention network integrating GPR and IRT data for automated bridge deck delamination detection with integrated uncertainty quantification. Our architecture combines three hierarchical attention mechanisms: temporal attention processing GPR signal sequences, spatial attention extracting thermal features from IRT imagery, and multi-head cross-modal attention with learnable modality embeddings discovering complementary defect patterns across sensors. We integrate comprehensive uncertainty quantification decomposing predictive uncertainty into epistemic and aleatoric components through Monte Carlo dropout and learned variance estimation, enabling selective prediction strategies that achieve 93% accuracy by routing uncertain cases to human review.
@@
Systematic experimental investigation across five bridge datasets from the SDNET2021 benchmark [8] reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five evaluated bridges. Under extreme imbalance exceeding 9:1 with minority classes below 2% representation, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become vulnerabilities. Comprehensive ablation studies demonstrate that cross-modal attention provides critical performance gains beyond within-modality attention alone, with multi-head mechanisms achieving substantially improved probabilistic calibration over single-head alternatives.
@@
Our cross-modal attention modules learn optimal feature interaction patterns between electromagnetic and thermal signatures, while Bayesian uncertainty quantification provides confidence intervals enabling risk-informed maintenance decisions [6]. On four datasets representing typical inspection scenarios, our architecture achieves 10–25 percentage point improvements in accuracy and AUC compared to single-modal approaches and concatenation-based fusion, with particularly strong results on the Park River Median dataset (96.6% accuracy, 99.8% AUC). Our investigation also characterizes boundary conditions where architectural simplification or specialized imbalance handling may prove preferable, providing the nuanced understanding essential for reliable deployment. Figure 1 provides an architectural overview of our proposed framework.
[…]
Figure 1: Architectural overview of the proposed multi-sensor attention network for automated bridge deck delamination detection. The framework processes Ground Penetrating Radar signals through temporal attention and Infrared Thermography images through spatial attention, integrating both modalities via cross-modal attention with learnable embeddings.
The remainder of this paper proceeds as follows. Section 2 surveys relevant literature on machine learning approaches for GPR-based and IRT-based bridge deck inspection, multi-modal NDE fusion methodologies, attention mechanisms for infrastructure defect detection, and uncertainty quantification techniques. Section 3 details our proposed multi-modal attention network architecture, including the temporal attention mechanism for GPR signal processing, spatial attention for IRT imagery analysis, cross-modal fusion with learnable modality embeddings, and the uncertainty quantification framework. Section 4 describes our experimental methodology using the SDNET2021 benchmark dataset and presents comprehensive results including hyperparameter optimization studies, ablation experiments, attention head configuration analysis, and per-bridge performance evaluation that characterizes both the method’s capabilities and its boundary conditions. Finally, Section 5 synthesizes our findings and identifies directions for future investigation. To facilitate reproducibility and enable further research, we make our implementation and trained models publicly available. [^1] Code and models available at: https://github.com/amoayedikia/mm-bridge-delam
@@ 2  Literature Review
Subsurface delamination in reinforced concrete bridge decks represents a critical structural concern that escapes detection through conventional visual inspection methods. The inability to identify chloride contamination, early-stage deterioration, and subsurface anomalies has driven extensive research into non-destructive evaluation (NDE) technologies, particularly Ground Penetrating Radar (GPR) and Infrared Thermography (IRT) [1]. While these sensing modalities offer complementary capabilities for bridge deck assessment, each exhibits fundamental limitations that have motivated development of multi-modal fusion strategies.
2.1  Machine Learning Approaches for GPR-Based Bridge Deck Assessment
Automated interpretation of GPR data through machine learning has transformed bridge deck inspection capabilities. Early comparative investigations established fundamental performance baselines for different algorithmic approaches. Ahmadvand et al. [17] evaluated 1D Convolutional Neural Networks against traditional machine learning methods including Support Vector Machines and K-Nearest Neighbors for classifying GPR A-scans from laboratory bridge deck specimens. Their findings demonstrated 1D-CNN superiority with 84% average accuracy and notably consistent performance (coefficient of variation: 0.04), establishing CNNs as the preferred approach for GPR-based delamination detection over conventional machine learning techniques.
Subsequent research explored hybrid frameworks combining signal processing with deep learning. Yu et al. [19] introduced a methodology integrating Synchrosqueezed Wavelet Transform for time-frequency feature extraction with VGG16 transfer learning, optimized through Modified Whale Optimization for hyperparameter selection. When applied to operational bridge deck data from North Dakota DOT, this hybrid approach exceeded 94% prediction accuracy, demonstrating that time-frequency decomposition effectively captures characteristics indicative of delamination signatures in electromagnetic reflection patterns.
Recent work has established important baselines on standardized benchmarks. Elseicy et al. [18] applied 1D-CNNs to the SDNET2021 benchmark dataset—the same dataset employed in the present investigation—achieving a weighted average F1-score of 0.994 when combining individual A-scan data with contextual information from neighboring signals. Their transfer learning experiments demonstrated robust generalization, achieving 92.6% accuracy when fine-tuned using only 5% of labels from new bridge data, establishing important performance benchmarks for GPR-based classification on this standardized dataset.
2.2  Deep Learning for IRT-Based Bridge Deck Delamination Detection
Thermal imagery analysis has progressed from traditional threshold-based techniques to sophisticated deep learning architectures capable of pixel-level delamination profiling. Cheng et al. [21] developed a custom encoder-decoder semantic segmentation architecture for analyzing UAV-based passive thermography of bridge decks. Their framework addressed three fundamental challenges in IRT analysis: variable delamination geometry and depth, environmental condition variations, and surface texture interference. Through strategic data augmentation, their approach improved intersection-over-union scores from 0.158 to 0.369, with successful generalization from controlled laboratory specimens to real-world bridge deck conditions.
More recently, Aljagoub et al. [20] advanced IRT-based detection by addressing training data scarcity through integration of real thermal images with numerically simulated thermal imagery across 39 distinct simulation scenarios. Their comparative evaluation of Mask R-CNN for instance segmentation versus YOLOv5 for object detection on UAV-based thermography from operational bridge decks successfully eliminated the subjective threshold requirements inherent in traditional IRT analysis while demonstrating robust delamination detection across varying defect properties. This work represents current state-of-the-art for deep learning-based thermal imagery analysis in bridge deck inspection.
2.3  Fusion Strategies for Multi-Modal NDE in Bridge Assessment
The complementary sensing capabilities of different NDE technologies have motivated fusion methodologies that integrate multiple sensing modalities for comprehensive bridge deck characterization. Mohamadi et al. [23] developed a feature-level fusion framework capable of combining up to eight NDE techniques—including GPR, Impact Echo, IRT, and Ultrasonic methods—employing Discrete Wavelet Transform for feature extraction followed by SVM classification. Laboratory bridge deck specimen experiments revealed that combining GPR with Impact Echo improved delamination detection accuracy by approximately 50% relative to individual techniques, with feature-level fusion consistently outperforming decision-level fusion strategies.
Neural network approaches have proven effective for learning cross-modal correlations in NDE data. Pashoutani et al. [24] introduced autoencoder architectures to model relationships between Vertical Electrical Impedance and GPR measurements acquired at spatially registered positions on reinforced concrete bridge decks. Field validation on a 55-meter operational steel-girder bridge demonstrated that autoencoder-based fusion generated condition maps with enhanced reliability and spatial resolution compared to individual NDE methods, establishing neural architectures as viable tools for cross-modal correlation learning in bridge deck NDE.
The SDNET2021 dataset introduction marked a significant advancement for multi-modal bridge deck inspection research, providing 663,102 annotated GPR signals, 1,936 Impact Echo signals, and 4,580,680 IRT pixels from five operational bridge decks [8]. The dataset’s three-class annotation scheme—intact, shallow delamination above reinforcement, and deep delamination below reinforcement—reflects standard bridge deck repair protocols and facilitates systematic evaluation of multi-modal fusion approaches across bridges exhibiting diverse structural characteristics.
2.4  Attention Mechanisms for Infrastructure Monitoring Applications
Attention mechanisms have emerged as transformative tools for infrastructure assessment applications, enabling context-dependent selective feature weighting. Temporal attention mechanisms have proven particularly effective for time-series sensor data processing. Chen et al. [25] developed a Temporal-Spatial Cross-Attention Network combining LSTM with Transformer architectures, achieving 92.6% accuracy on subsurface defect detection while reducing computational demands by 40% relative to baseline approaches. This demonstrated that temporal attention effectively identifies critical reflection patterns in sequential GPR data.
Spatial attention mechanisms have shown substantial performance improvements for thermal imagery analysis. Wang et al. [26] demonstrated that combining channel and spatial attention mechanisms yields 10–12% AUC improvements over baseline CNN approaches for thermal defect detection, establishing spatial attention as an effective architectural component for infrastructure applications involving thermal imagery.
Cross-modal attention represents the current frontier in multi-sensor fusion, enabling learning of inter-modality relationships that capture complementary defect signatures across sensing modalities. Lang and Lv [10] demonstrated that attention-based fusion mechanisms outperform traditional concatenation strategies by 10–30% across diverse infrastructure applications by modeling sophisticated correspondences between different sensing modalities. Vision Transformer architectures adapted for infrastructure defect detection have achieved accuracy rates exceeding 90% while outperforming CNN-based approaches [9]. Dynamic sparse attention architectures have further advanced computational efficiency, reducing complexity by 60% while maintaining detection accuracy [11].
2.5  Uncertainty Quantification for Safety-Critical Infrastructure Assessment
Safety-critical infrastructure applications demand not only accurate predictions but also well-calibrated confidence estimates for decision-making. Monte Carlo Dropout has emerged as the predominant approach for uncertainty quantification in deep learning, with empirical research demonstrating that 15–25 forward passes provide reliable epistemic uncertainty estimates [6].
Pantoja-Rosero et al. [29] advanced uncertainty-aware defect detection through their Bayesian Boundary-Aware Convolutional Network, achieving 84.7% F1-score while providing decomposed uncertainty estimates that separate model uncertainty from data uncertainty. Their analysis revealed strong correlations between mean class softmax variance, entropy, and misclassification events, enabling uncertainty-triggered human intervention protocols for ambiguous detection cases.
2.6  Class Imbalance Challenges in Infrastructure Defect Datasets
Severe class imbalance inherent in bridge deck inspection data—exemplified by SDNET2021’s typical distribution of 77% intact regions, 17% shallow delamination, and 6% deep delamination—has motivated specialized training methodologies. Jadon [28] conducted comprehensive evaluation demonstrating that Focal Tversky Loss outperformed 11 alternative loss functions across infrastructure benchmarks, achieving 23% improvement in minority class detection compared to standard cross-entropy. Rahman and Wang [30] showed that weighted cross-entropy with inverse frequency weighting provides comparable performance with simpler implementation requirements.
However, existing class imbalance research has focused predominantly on loss function modifications, with limited investigation of how architectural choices interact with class imbalance. Specifically, attention mechanisms—with their increased parameter capacity and learned weighting schemes—may exhibit different sensitivity to class imbalance compared to simpler architectures, yet this relationship remains poorly characterized in the literature.
2.7  Identified Research Gaps and Study Objectives
Despite substantial progress in both GPR-based and IRT-based machine learning for bridge deck inspection, several critical knowledge gaps persist. First, while existing multi-modal fusion approaches for bridge decks have employed traditional machine learning (SVM, autoencoders), attention-based deep learning architectures that could model complex inter-modal relationships between electromagnetic and thermal signatures remain unexplored for this application. Second, although 1D-CNNs have proven effective for GPR processing and encoder-decoder networks for IRT segmentation, no prior work has integrated temporal attention for GPR signals with spatial attention for IRT imagery within a unified cross-modal framework specifically for bridge deck assessment.
Third, uncertainty quantification research has developed largely independently from attention mechanism research in bridge deck inspection contexts, with limited investigation of whether attention-based architectures produce better-calibrated confidence estimates—a critical consideration for safety-critical infrastructure applications. Finally, the five bridge datasets within SDNET2021 exhibit substantially different class distributions (ranging from relatively balanced to severely imbalanced), yet prior work using this benchmark has not systematically investigated how model performance varies with dataset characteristics. This leaves practitioners without evidence-based guidance for selecting appropriate methods for specific inspection scenarios encountered in practice.
This work addresses these gaps through introduction of a multi-modal attention network integrating GPR temporal patterns with IRT spatial signatures via cross-modal attention mechanisms, coupled with comprehensive uncertainty quantification. Rather than claiming universal architectural superiority, we systematically characterize the conditions under which attention-based fusion provides advantages over simpler alternatives, providing practitioners with actionable deployment guidance for automated inspection systems across real-world bridge infrastructure networks.